# COMAE: COMprehensive Attribute Exploration for Zero-shot Hashing

### v1

@@
Abstract
@@
keywords:
[…]
1 Introduction
@@
Image hashing techniques have garnered substantial interest due to their remarkable efficacy in multimedia retrieval and content-based searches for images and documents [18, 41, 16]. Particularly, hashing techniques convert high-dimensional and complicated data into binary codes, which enables efficient searches in the binary Hamming space, thereby contributing to a smaller memory footprint and increased search efficiency [32, 20, 15]. The capability to depict intricate data in a compact binary format not only accelerates retrieval processes but also markedly improves storage efficiency. Consequently, image hashing emerges as a highly valuable asset in situations that necessitate swift and resource-efficient information retrieval.
@@
With the popularity of content and social apps, the growth of Internet images leads to sort of new concepts, presenting a challenge for traditional hashing algorithms that struggle to adapt to unseen data in previously unobserved categories [38, 41, 23]. Recent researchers paid attention to the Zero-Shot Hashing (ZSH) issue. Visual attributes, pervasive across prevalent image datasets, serve to delineate discriminative visual properties intrinsic to objects, transcending specific classes and establishing a foundation for addressing the intricacies of previously unseen data [17, 4]. Notably, these attributes offer a robust representation for recognizing out-of-distribution samples, making them valuable for capturing similarities in the seen and unseen concepts in Zero-Shot Learning (ZSL) tasks [2, 45]. Fig. 1 shows an example of the class-level attributes. Nevertheless, the exploration of visual attributes within the realm of ZSH tasks has been relatively meager. AH [45] adopts a straightforward approach by constructing a multi-task architecture aimed at comprehensively capturing the intricate relationships between visual attributes and category classes, thereby facilitating the transfer of knowledge.
Figure 1: An example of the class-level attributes.
@@
While considerable success has been achieved in utilizing visual attributes to establish connections among visual attributes, image features, class labels, and binary hash codes, there still exists urgent limitations [20, 45]. Firstly, image locality representations are proven to be critical to zero-shot learning tasks [19], as the locality exhibits more effective transferability between seen classes and unseen classes. However, existing works ignore the locality representations in deep hashing tasks, which decreases the upper bound of the transferability in hashing algorithms. Secondly, existing works mainly focus on individual image representation learning, optimizing each instance independently. Each dimension of the visual attribute is a continuous value. In contrast to the binary one-hot class encoding, these continuous values encompass richer and more nuanced distance-related information. These informative attributes are aptly poised to capture intricate pairwise relationships, while they are largely ignored by prevailing works. Thirdly, existing works generally employ a multi-task learning framework to independently utilize attribute labels and class labels. Yet their actual relationships are more complicated than such simple presumes. To be specific, two images may belong to the same class but pose disparate attributes. For example, two images with a “bird” class may have “black leg” and “yellow leg”. Similarly, instances with different classes may also exhibit sharing analogous attributes. For example, two images with “bird” and “wapiti” classes may have the same “black leg”. The existing methods may not lead to the global optimum, thus necessitating performing joint optimization to achieve the overall optimum for both tasks.
@@
2 Related Work
2.1 Deep Hashing
@@
Image hashing is a popularly used method for approximating nearest neighbor search in large-scale retrieval scenarios [38]. Shallow image hashing methods often leverage handcrafted features or employ traditional machine-learning algorithms to generate compact binary representations. Typical works include LSH [1] and ITQ [10]. Shallow methods are computationally efficient and exhibit simplicity in their design, yet they may struggle to capture complex and high-level semantic features inherent in images.
@@
Over the past decades, deep learning models have shown excellent progress in the field of Computer Vision [17, 27] and Natural Language Processing [33, 23], many scholars pay attention to leveraging deep neural networks to learn the abstract representations. Deep hashing methods can be achieved through three paths. The first path utilizes limited information, which has a lower upper bound of the capability. ASZH [32] is a typical work in the first path, that learns hash functions based on the seeable category labels, for generating binary codes for unseeable data. The second path [49, 11] generates semantic vectors based on the visual attributes to find a semantic embedding space, then transfers the common knowledge to unseeable classes. TSK, SitNet, and OPZH [46, 11, 49] are representative semantic-based works, they conduct a multi-task framework in which the class labels and attribute vectors are co-trained through the hashing network, and the labels are projected into a semantic embedding space. However, there exists an inherent gap between the semantic and visual space, and the alignment of spaces is tough. The third path [45] constructs relationships among attributes, image features, binary codes, and class labels. These models are considered to be more generalized when dealing with out-of-distribution and open-set samples [2, 9].
2.2 Zero-shot Learning
@@
Zero-shot learning (ZSL) [8, 44, 51, 22] is designed to extend object recognition across known and unknown classes via a unified embedding space, where both categories are described by their semantic visual identifiers. Early ZSL methods [8, 27] harvest global visual features from networks, whether pre-trained or capable of end-to-end training. Typically, end-to-end frameworks often surpass pre-trained ones by refining visual attributes, thus mitigating biases between the datasets used for training and those for ZSL applications. Nonetheless,such strategies typically produce less than ideal outcomes [44], as they struggle to discern the subtle variances between seen and unseen categories. Recent innovations in ZSL predominantly concentrate on methods based on attention [51, 36, 44], employing visual attributes to identify distinct feature regions, for instance, specific parts of a bird. This emphasis highlights the critical roles of specificity and the arrangement of image features in ZSL. Proposals for prototype networks [36] strive to pinpoint various attributes within images to bolster the precision of feature localization. Inspired by these insights into visual attributes and the delicate distinctions between recognized and unrecognized classes, our methodology adopts the prototype network to intricately map the relationships between visual attributes and hash codes [44].
@@
Figure 2: The architecture of the proposed COMAE. It consists of three modules: a) Point-wise objective aims at improving image locality and attribute representations; b) Pair-wise loss is proposed to learn the representations from individual learning to context-based learning; and c) Class-wise constraint is designed to capture relationships of attributes and class labels.
2.3 Contrastive Learning
@@
Contrastive learning aims at honing representations by juxtaposing similar instance pairs (positive pairs) against dissimilar ones (negative pairs) [12, 13], enhancing the model’s ability to differentiate between closely related and distinct instances effectively. Many works demonstrate that contrastive learning is able to alleviate the problems of numerous categories and imbalanced class distributions. In addition, it can amplify the subtle distinctions among similar classes, which assists in capturing intricate details that are often overlooked by basic models. [12] is the first attempt work to improve learned representations by contrasting positive pairs and negative pairs. SimCLR [5] utilizes a straightforward self-supervised learning architecture that substitutes the conventional memory bank with batch-specific elements, significantly enhancing its performance on the ImageNet dataset. As hash code is a form of representation, recent researches [47, 34, 28] take attempts to utilize contrastive learning in deep hashing. [47] proposes weighted contrastive hashing to explore the fine-grained semantic relations among images. [34] propose to However, random sampling could result in augmented images with different semantics, deteriorating the performance of contrastive learning.
3 Method
@@ 3.1 Problem Definition
3.2 Point-wise Attribute Consistency
Image locality representations are proven to be practical in transferring knowledge from seen to unseen classes [44, 19]. For more subtly depicted image and attribute characterizations, we first conduct point-wise consistency constraint, which takes the attribute regression loss as the supervision. Prototype learning has proven to be robust in dealing with open-set recognition and out-of-distribution samples [44, 37] in recent years. In the prototype learning scheme, each category is represented by a prototype. This prototype can be the mean of the sample in the category, the center of mass, or some other representative sample. Thus we introduce a prototype module aiming at capturing the relationships of local features and visual attributes.
@@ 3.2 Point-wise Attribute Consistency
$M_{i,j}^{k}$ $=\Theta\left(p_{k},f_{i,j}(x)\right),$ (1)
@@
$\mathcal{L}_{\text{pointwise}}=\sum_{i}^{N}\frac{1}{N}||a_{i}-\hat{a_{i}}||^{2}.$ (2)
Furthermore, we refine the similarity map to intensively focus on its peak region. To achieve this, we employ a compactness regularizer [50] on each similarity map $M^{k}$, ensuring a concentrated similarity assessment. The equation is defined as,
$\mathcal{L}_{\text{reg}}=\sum^{K}_{k}\sum^{H}_{h}\sum^{W}_{w}M_{i,j}^{k}[(i-\tilde{i})^{2}+(j-\tilde{j})^{2}],$ (3)
@@
3.3 Pair-wise Attribute Consistency
To make full use of the limited yet crucial pieces of information, we conduct pair-wise attribute exploration to depict the context relationships of continuous-valued attributes. Recent studies proved that contrastive learning can enhance the discriminative capability of feature representations and enable the establishment of meaningful relationships between seen and unseen classes [13, 28, 6]. The fundamental concept of pairwise loss revolves around learning feature representations by comparing the similarity between different pairs of samples. The pairwise loss function enhances the model’s ability to compare sample pairs across these groups, directing it to recognize that samples within the same category should exhibit higher similarity compared to those from different categories. This approach ensures a more discerning and category-aware similarity assessment within the model’s learning process. Thus we utilize contrastive learning in the pair-wise constraint. We begin with constructing the positive sample set $S^{+}_{d}(i)$ of the $i$-th image and the negative sample set $S^{-}_{d}(i)$ in the $d$-th dimension of all attributes. Specifically, given the attribute of $i$-th image, which denoted as $a_{i}$, we build the positive set of $i$-th image as follows,
$S^{+}_{d}(i)=\{j\mid\lvert a_{i,d}-a_{j,d}\rvert<\epsilon,j\neq i,j=1,...,N\},$ (4)
@@ 3.3 Pair-wise Attribute Consistency
$\mathcal{L}_{\text{pairwise}}=-\sum_{d}^{D}\sum_{i}^{N}\frac{1}{DN}\log\frac{\sum_{j\in S^{+}_{d}(i)}\text{exp}(\hat{a_{i}}\hat{a_{j}}/\tau)}{\sum_{t\in S_{d}(i)}\text{exp}(\hat{a_{i}}\hat{a_{t}}/\tau)},$ (5)
@@
Algorithm 1 Algorithm of COMAE
1:Training Images $X=\{x_{i}\}_{i=1}^{N}$, class labels $Y=\{y_{i}\}_{i=1}^{N}$, and image attribute vectors $A=\{a_{i}\}_{i=1}^{N}$.
2:The parameters $\Theta$
3:for $i=0,...,N$ do
4:     Generate image initial embeddings $f(x_{i})$ based on the image encoder, i.e., ResNet101.
5:Generate the predicted prototypes $\{M^{k}\}_{k=1}^{K}$ for each attribute.
6:for $i=0,...,N$ do
7:     Construct the positive sample set $S^{+}_{d}(i)$ of the $i$-th image and negative sample set $S^{-}_{d}(i)$, as Eqn. (4).
8:while not converge do :
9:     Compute the pointwise loss, as Eqn. (2).
10:     Compute the pairwise loss, as Eqn. (5).
11:     Compute the classwise loss, as Eqn. (6).
12:     for $i=0,...,N$ do
13:         Generate binary hashing codes $B=\{b_{i}\}_{i=1}^{N}$ based on image representations $f(x_{i})$ through Eqn. (7).
14:     Compute the joint loss through Eqn. (9)
15:     Compute the backpropagation error.
16:     Update the parameters $\Theta$ and image hidden representations $f(x_{i})$. return Image hashing codes $B=\{b_{i}\}_{i=1}^{N},b_{i}\in\{-1,1\}^{l}$.
3.4 Class-wise Attribute Exploration
@@
$\mathcal{L}_{\text{ce}}=-\frac{1}{|Y^{s}|}\sum_{y_{i}\in Y^{s}}\log\frac{\text{exp}\left(g(x_{i})^{T}Wa_{i}\right)}{\text{exp}\left(g(x_{i})^{T}W\hat{a_{i}}\right)},$ (6)
@@ 3.5 Hashing Learning
$b_{i}=\text{sign}\left(\sum_{c=1}^{C}\sum_{h=1}^{H}\sum_{w=1}^{W}f(x_{i})W_{h,w,c}\right),$ (7)
@@
To compress intra-class hashing code with strong constraints, we use the hypersphere loss [21, 40] to optimize the FCH layer with large angular boundaries between classes. The spherical hashing objective offers several advantages within the context of similarity-preserving hash coding. One notable advantage lies in its ability to enforce constraints on binary codes such that they reside on a hypersphere in the Hamming space. This geometric constraint enhances the discriminative power of the learned hash codes by facilitating a more structured and separable representation of data points. Additionally, the hypersphere constraint promotes semantic similarity preservation, as points closer to the hypersphere correspond to more semantically similar items. This results in improved retrieval performance, particularly in scenarios where maintaining semantic relationships among binary codes is crucial, such as image retrieval and nearest neighbor search. Furthermore, the geometric regularity induced by the hypersphere constraint contributes to robustness against noise and variations, thereby enhancing the overall effectiveness and stability of the hashing model. It is defined as,
$\mathcal{L}_{\text{hash}}$ $=-\frac{1}{N}\sum_{i}^{N}\log\left(\frac{e^{||b_{i}||\psi(\theta_{yi,i})}}{e^{||b_{i}||\psi(\theta_{yi,i})}+\sum_{j\neq y_{i}}e^{||b_{i}||\cos(\theta_{j},i)}}\right),$ (8)
@@
$\mathcal{L}=\sum\lambda_{i}\mathcal{L}_{i},$ (9)
@@
The pseudocode of COMAE is shown in Algorithm. 1.
Table 1: The mAP@5000 results on three popular ZSH datasets in different code lengths.

### v2

@@
Abstract.
Zero-shot hashing (ZSH) has shown excellent success owing to its efficiency and generalization in large-scale retrieval scenarios. While considerable success has been achieved, there still exist urgent limitations. Existing works ignore the locality relationships of representations and attributes, which have effective transferability between seeable classes and unseeable classes. Also, the continuous-value attributes are not fully harnessed. In response, we conduct a COMprehensive Attribute Exploration for ZSH, named COMAE, which depicts the relationships from seen classes to unseen ones through three meticulously designed explorations, i.e., point-wise, pair-wise and class-wise consistency constraints. By regressing attributes from the proposed attribute prototype network, COMAE learns the local features that are relevant to the visual attributes. Then COMAE utilizes contrastive learning to comprehensively depict the context of attributes, rather than instance-independent optimization. Finally, the class-wise constraint is designed to cohesively learn the hash code, image representation, and visual attributes more effectively. Experimental results on the popular ZSH datasets demonstrate that COMAE outperforms state-of-the-art hashing techniques, especially in scenarios with a larger number of unseen label classes. Our codes are available at https://github.com/itsnotacie/TOMM2024-COMAE.
Abstract.
@@
Prototype Learning, Local Representation, Zero-shot Hashing, Image Zero-shot Hashing, Image Hashing, Image Retrieval.
copyright: acmlicensedjournalyear: 2024doi: XXXXXXjournal: TOMMjournalvolume: 1journalnumber: 1article: 1publicationmonth: 1ccs: Computing methodologies Visual content-based indexing and retrievalccs: Information systems Image search
@@
Image hashing techniques have garnered substantial interest due to their remarkable efficacy in multimedia retrieval and content-based searches for images and documents (Lew et al., 2006; Wang et al., 2017; Ju et al., 2024). Particularly, hashing techniques convert high-dimensional and complicated data into binary codes, which enables efficient searches in the binary Hamming space, thereby contributing to a smaller memory footprint and increased search efficiency (Shi et al., 2022; Li and van Gemert, 2021; Ju et al., 2023). The capability to depict intricate data in a compact binary format not only accelerates retrieval processes but also markedly improves storage efficiency. Consequently, image hashing emerges as a highly valuable asset in situations that necessitate swift and resource-efficient information retrieval.
@@
With the popularity of content and social apps, the growth of Internet images leads to sort of new concepts, presenting a challenge for traditional hashing algorithms that struggle to adapt to unseen data in previously unobserved categories (Venkatesan et al., 2000; Wang et al., 2017; Long et al., 2021). Recent researchers paid attention to the Zero-Shot Hashing (ZSH) issue. Visual attributes, pervasive across prevalent image datasets, serve to delineate discriminative visual properties intrinsic to objects, transcending specific classes and establishing a foundation for addressing the intricacies of previously unseen data (Khan et al., 2018; Chen et al., 2022). Notably, these attributes offer a robust representation for recognizing out-of-distribution samples, making them valuable for capturing similarities in the seen and unseen concepts in Zero-Shot Learning (ZSL) tasks (Arik and Pfister, 2019; Xu et al., 2017). Fig. 1 shows an example of the class-level attributes. Nevertheless, the exploration of visual attributes within the realm of ZSH tasks has been relatively meager. AH (Xu et al., 2017) adopts a straightforward approach by constructing a multi-task architecture to capture the intricate relationships between visual attributes and category classes, thereby facilitating the transfer of knowledge.
@@
While considerable success has been achieved in utilizing visual attributes to establish connections among visual attributes, image features, class labels, and binary hash codes, there still exists urgent limitations (Li and van Gemert, 2021; Xu et al., 2017). Firstly, image locality representations are proven to be critical to zero-shot learning tasks (Li et al., 2021), as the locality exhibits more effective transferability between seen classes and unseen classes. However, existing works ignore the locality representations in deep hashing tasks, which decreases the upper bound of the transferability in hashing algorithms. Secondly, existing works mainly focus on individual image representation learning, optimizing each instance independently. Each dimension of the visual attribute is a continuous value. In contrast to the binary one-hot class encoding, these continuous values encompass richer and more nuanced distance-related information. These informative attributes are aptly poised to capture intricate pairwise relationships, while they are largely ignored by prevailing works. Thirdly, existing works generally employ a multi-task learning framework to independently utilize attribute labels and class labels. Yet their actual relationships are more complicated than such simple presumes. To be specific, two images may belong to the same class but pose disparate attributes. For example, two images with a “bird” class may have “black leg” and “yellow leg”. Similarly, instances with different classes may also exhibit sharing analogous attributes. For example, two images with “bird” and “wapiti” classes may have the same “black leg”. The existing methods may not lead to the global optimum, thus necessitating performing joint optimization to achieve the overall optimum for both tasks.
@@
1. Related Work
1.1. Deep Hashing
@@
Image hashing is a popularly used method for approximating nearest neighbor search in large-scale retrieval scenarios (Venkatesan et al., 2000). Shallow image hashing methods often leverage handcrafted features or employ traditional machine-learning algorithms to generate compact binary representations. Typical works include LSH (Andoni and Indyk, 2008) and ITQ (Gong et al., 2012). Shallow methods are computationally efficient and exhibit simplicity in their design, yet they may struggle to capture complex and high-level semantic features inherent in images.
@@
Over the past decades, deep learning models have shown excellent progress in the field of Computer Vision (Khan et al., 2018; Pourpanah et al., 2022) and Natural Language Processing (Song et al., 2020; Long et al., 2021), many scholars pay attention to leveraging deep neural networks to learn the abstract representations. Deep hashing methods can be achieved through three paths. The first path utilizes limited information, which has a lower upper bound of the capability. ASZH (Shi et al., 2022) is a typical work in the first path, that learns hash functions based on the seeable category labels, for generating binary codes for unseeable data. The second path (Zhang et al., 2019; Guo et al., 2017) generates semantic vectors based on the visual attributes to find a semantic embedding space, then transfers the common knowledge to unseeable classes. TSK, SitNet, and OPZH (Yang et al., 2016; Guo et al., 2017; Zhang et al., 2019) are representative semantic-based works, they conduct a multi-task framework in which the class labels and attribute vectors are co-trained through the hashing network, and the labels are projected into a semantic embedding space. However, there exists an inherent gap between the semantic and visual space, and the alignment of spaces is tough. The third path (Xu et al., 2017) constructs relationships among attributes, image features, binary codes, and class labels. These models are considered to be more generalized when dealing with out-of-distribution and open-set samples (Arik and Pfister, 2019; Gao et al., 2019).
1.2. Zero-shot Learning
@@
Zero-shot learning (ZSL) (Farhadi et al., 2009; Xu et al., 2020; Zheng and Gupta, 2022; Long et al., 2020) is designed to extend object recognition across known and unknown classes via a unified embedding space, where both categories are described by their semantic visual identifiers. Early ZSL methods (Farhadi et al., 2009; Pourpanah et al., 2022) harvest global visual features from networks, whether pre-trained or capable of end-to-end training. Typically, end-to-end frameworks often surpass pre-trained ones by refining visual attributes, thus mitigating biases between the datasets used for training and those for ZSL applications. Nonetheless,such strategies typically produce less than ideal outcomes (Xu et al., 2020), as they struggle to discern the subtle variances between seen and unseen categories. Recent innovations in ZSL predominantly concentrate on methods based on attention (Zheng and Gupta, 2022; Sylvain et al., 2019; Xu et al., 2020), employing visual attributes to identify distinct feature regions, for instance, specific parts of a bird. This emphasis highlights the critical roles of specificity and the arrangement of image features in ZSL. Proposals for prototype networks (Sylvain et al., 2019) strive to pinpoint various attributes within images to bolster the precision of feature localization. Inspired by these insights into visual attributes and the delicate distinctions between recognized and unrecognized classes, our methodology adopts the prototype network to intricately map the relationships between visual attributes and hash codes (Xu et al., 2020).
@@
Figure 2. The architecture of the proposed COMAE. It consists of three modules: a) Point-wise objective aims at improving image locality and attribute representations; b) Pair-wise loss is proposed to learn the representations from individual learning to context-based learning; and c) Class-wise constraint is designed to capture relationships of attributes and class labels.
1.3. Contrastive Learning
@@
Contrastive learning aims at honing representations by juxtaposing similar instance pairs (positive pairs) against dissimilar ones (negative pairs) (Hadsell et al., 2006; Han et al., 2021), enhancing the model’s ability to differentiate between closely related and distinct instances effectively. Many works demonstrate that contrastive learning is able to alleviate the problems of numerous categories and imbalanced class distributions. In addition, it can amplify the subtle distinctions among similar classes, which assists in capturing intricate details that are often overlooked by basic models. (Hadsell et al., 2006) is the first attempt work to improve learned representations by contrasting positive and negative pairs. SimCLR (Chen et al., 2020) utilizes a self-supervised learning architecture that substitutes the conventional memory bank with batch-specific elements, significantly enhancing its performance on the ImageNet dataset. As hash code is a form of representation, recent researches (Yu et al., 2022; Song et al., 2023; Qiu et al., 2021) take attempts to utilize contrastive learning in deep hashing. (Yu et al., 2022) proposes weighted contrastive hashing to explore the fine-grained semantic relations among images. (Song et al., 2023) propose to However, random sampling could result in augmented images with different semantics, deteriorating the performance of contrastive learning.
2. Method
@@ 2.1. Problem Definition
2.2. Point-wise Attribute Consistency
Image locality representations are proven to be practical in transferring knowledge from seen to unseen classes (Xu et al., 2020; Li et al., 2021). For more subtly depicted image and attribute characterizations, we first conduct point-wise consistency constraint, which takes the attribute regression loss as the supervision. Prototype learning has proven to be robust in dealing with open-set recognition and out-of-distribution samples (Xu et al., 2020; Tan et al., 2022) in recent years. In the prototype learning scheme, each category is represented by a prototype. This prototype can be the mean of the sample in the category, the center of mass, or some other representative sample. Thus we introduce a prototype module aiming at capturing the relationships of local features and visual attributes.
@@ 2.2. Point-wise Attribute Consistency
(1) $M_{i,j}^{k}$ $=\Theta\left(p_{k},f_{i,j}(x)\right),$
@@
(2) $\mathcal{L}_{\text{pointwise}}=\sum_{i}^{N}\frac{1}{N}||a_{i}-\hat{a_{i}}||^{2}.$
Furthermore, we refine the similarity map to intensively focus on its peak region. To achieve this, we employ a compactness regularizer (Zheng et al., 2017) on each similarity map $M^{k}$, ensuring a concentrated similarity assessment. The equation is defined as,
(3) $\mathcal{L}_{\text{reg}}=\sum^{K}_{k}\sum^{H}_{h}\sum^{W}_{w}M_{i,j}^{k}[(i-\tilde{i})^{2}+(j-\tilde{j})^{2}],$
@@
2.3. Pair-wise Attribute Consistency
To make full use of the limited yet crucial pieces of information, we conduct pair-wise attribute exploration to depict the context relationships of continuous-valued attributes. Recent studies proved that contrastive learning can enhance the discriminative capability of feature representations and enable the establishment of meaningful relationships between seen and unseen classes (Han et al., 2021; Qiu et al., 2021; Chen et al., 2018). The fundamental concept of pairwise loss revolves around learning feature representations by comparing the similarity between different pairs of samples. The pairwise loss function enhances the model’s ability to compare sample pairs across these groups, directing it to recognize that samples within the same category should exhibit higher similarity compared to those from different categories. This approach ensures a more discerning and category-aware similarity assessment within the model’s learning process. Thus we utilize contrastive learning in the pair-wise constraint. We begin with constructing the positive sample set $S^{+}_{d}(i)$ of the $i$-th image and the negative sample set $S^{-}_{d}(i)$ in the $d$-th dimension of all attributes. Specifically, given the attribute of $i$-th image, which denoted as $a_{i}$, we build the positive set of $i$-th image as follows,
(4) $S^{+}_{d}(i)=\{j\mid\lvert a_{i,d}-a_{j,d}\rvert<\epsilon,j\neq i,j=1,...,N\},$
@@ 2.3. Pair-wise Attribute Consistency
(5) $\mathcal{L}_{\text{pairwise}}=-\sum_{d}^{D}\sum_{i}^{N}\frac{1}{DN}\log\frac{\sum_{j\in S^{+}_{d}(i)}\text{exp}(\hat{a_{i}}\hat{a_{j}}/\tau)}{\sum_{t\in S_{d}(i)}\text{exp}(\hat{a_{i}}\hat{a_{t}}/\tau)},$
@@
2.4. Class-wise Attribute Exploration
@@
(6) $\mathcal{L}_{\text{ce}}=-\frac{1}{|Y^{s}|}\sum_{y_{i}\in Y^{s}}\log\frac{\text{exp}\left(g(x_{i})^{T}Wa_{i}\right)}{\text{exp}\left(g(x_{i})^{T}W\hat{a_{i}}\right)},$
@@ 2.5. Hashing Learning
(7) $b_{i}=\text{sign}\left(\sum_{c=1}^{C}\sum_{h=1}^{H}\sum_{w=1}^{W}f(x_{i})W_{h,w,c}\right),$
@@
To compress intra-class hashing code with strong constraints, we use the hypersphere loss (Liu et al., 2017; Wang et al., 2018) to optimize the FCH layer with large angular boundaries between classes. The spherical hashing objective offers several advantages within the context of similarity-preserving hash coding. One notable advantage lies in its ability to enforce constraints on binary codes such that they reside on a hypersphere in the Hamming space. This geometric constraint enhances the discriminative power of the learned hash codes by facilitating a more structured and separable representation of data points. Additionally, the hypersphere constraint promotes semantic similarity preservation, as points closer to the hypersphere correspond to more semantically similar items. This results in improved retrieval performance, particularly in scenarios where maintaining semantic relationships among binary codes is crucial, such as image retrieval and nearest neighbor search. Furthermore, the geometric regularity induced by the hypersphere constraint contributes to robustness against noise and variations, thereby enhancing the overall effectiveness and stability of the hashing model. It is defined as,
(8) $\mathcal{L}_{\text{hash}}$ $=-\frac{1}{N}\sum_{i}^{N}\log\left(\frac{e^{||b_{i}||\psi(\theta_{yi,i})}}{e^{||b_{i}||\psi(\theta_{yi,i})}+\sum_{j\neq y_{i}}e^{||b_{i}||\cos(\theta_{j},i)}}\right),$
@@
(9) $\mathcal{L}=\sum\lambda_{i}\mathcal{L}_{i},$
@@
Table 1. The mAP@5000 results on three popular ZSH datasets in different code lengths.